OpenStudy (anonymous):

Anyone quickly remember how to break up (x^3 - 1)

OpenStudy (anonymous):

http://www.purplemath.com/modules/specfact2.htm a³ – b³ = (a – b)(a² + ab + b²) x³ - 1 = (x - 1)(x² + x + 1)

OpenStudy (anonymous):

alright cool thanks